# Exact Consistency Under Partial Views: Graph Colorability, Capacity, and Equality in Multi-Location Encodings

### 16 Mar 2026 v3

@@ Abstract
We construct a structural theory of failure and a graph-capacity framework for analyzing structural integrity under distributed source coding. Structural integrity means the code can correct erasures: the mapping from source to observable syndrome is injective. Admissible partial views induce a confusability graph on latent tuples; in the exact coordinate-view model, this graph class is exactly characterized by upward-closed families of coordinate-agreement sets, and exact recovery with a $T$-ary tag is equivalent to $T$-colorability. Repeated composition yields strong powers, so the normalized block-rate sequence converges to asymptotic Shannon capacity bounded above by Lovász-$\vartheta$. The upper theory is sharp whenever confusability is transitive; meet-witnessing and fiber coherence provide checkable sufficient conditions for that collapse. Under an affine restriction, the coordinate structure yields a representable matroid whose rank bounds confusability. The theory yields verifiable structural integrity criteria with applications to programming-language runtimes, databases, and dependency managers.
footnotetext: © 2026 Tristan Simas. This work is licensed under $C$ BY 4.0. License: https://creativecommons.org/licenses/by/4.0/
@@ I Introduction
The substantive question in distributed source coding with side information is what exact ambiguity structure survives once a system exposes only partial observations of its source, and what formal structural integrity guarantees are possible in that regime. Structural integrity here means the code can correct erasures: the mapping from source to observable syndrome is injective, so the latent state is uniquely recoverable from the observable tag. This paper answers that question with a zero-error graph-capacity theory for deterministic partial views on latent tuples.
The theory is predictive: given only the view-family specification, the full failure topology, exact recovery law, asymptotic capacity, and standard upper bounds are deductive consequences of the architecture, with the realizable graph class fully characterized.
When one latent fact or tuple of facts is represented at several modifiable locations, the available observation need not identify a unique latent state. Under partial views, the surviving ambiguity has structure: the admissible view family induces a confusability graph on latent tuples. Its edges record the exact state pairs that the architecture cannot separate, proper colorings quantify the side-information budget needed for exact recovery, and repeated composition pushes the same structure to strong powers and asymptotic Shannon capacity. The paper therefore studies not just whether a multi-location system can fail, but the topology of those failures, the integrity guarantees excluded or permitted by that topology, and the recovery laws forced by it. The resulting theory lies at the intersection of zero-error information theory, side-information source coding, finite converses, and structural integrity analysis [1, 2, 3, 4, 5, 6].
Novelty. Unlike Witsenhausen’s setting where the observation law is given, here a deterministic multi-location partial-view architecture generates that law and thereby constrains the realizable graph class.(L: MFT125-131)
The graph-generation mechanism is now more sharply characterized. On the full labeled tuple space, the exact coordinate-view model realizes exactly the confusability relations determined by upward-closed families of coordinate-agreement sets: two distinct tuples are adjacent if and only if their agreement set lies in such a family, and conversely every upward-closed family is realizable by some view architecture.(L: MFT125-128) When the alphabet has at least two symbols, every proper agreement set occurs for some distinct tuple pair, so this is a complete characterization up to the vacuous full-agreement case of identical tuples.(L: MFT129-131) Coordinatewise alphabet permutations still act by graph automorphisms, and any latent tuple can be carried to any other by such an automorphism.(L: MFT120-124) Hence the induced graphs form a specific monotone agreement-set class rather than an arbitrary finite graph family. The model already generates non-clique graphs, includes the cluster-collapse subclass on which equality is explicit, and is closed under the block-composition operation that becomes strong product on confusability graphs. The binary square already yields a $4$-cycle rather than a clique, and its iterates give a concrete infinite non-clique strong-product family inside the class.
@@ I-A Distributed Source Coding Formulation
Failure geometry from partial views. The main jump of the paper is from clique-shaped ambiguity to architecture-specific failure structure. In the multi-fact partial-observation extension, the confusability graph is the structural object controlling exact consistency under partial views, exact recovery is equivalent to graph colorability, and exact finite weighted success is determined by the maximum $T$-colorable induced subgraph. The binary two-fact square already shows that the induced graph need not be a clique: it is a $4$-cycle, so failures above the unit-rate boundary are structured rather than uniform. (L: MFT3, MFT5-10, MFT24)
@@
Asymptotic rate and upper theory. This graph characterization then lifts to strong powers and asymptotic zero-error rates. The full $n$-block confusability graph is the $n$-fold strong power of the one-shot graph; the normalized block-rate sequence converges to a real asymptotic Shannon-capacity value equal to the supremum of the finite block rates; and that value is bounded above by both the logarithm of the complement chromatic number and a fixed Lovász-$\vartheta$ upper convention. The one-shot upper object matches standard orthonormal-representation, primal-PSD, and dual-theta forms. (L: MFT35, MFT48, MFT55, MFT64, MFT69, MFT84)
@@
Equality characterization. The paper also proves an equality characterization. For the original clique-fiber subclass coming from a surjective label map, asymptotic Shannon capacity and the fixed Lovász-$\vartheta$ upper both collapse to $\log|\mathcal{B}|$, where $\mathcal{B}$ is the fiber-label alphabet. This equality for cluster-graph structure is classical; the new point here is that transitivity of base confusability gives a model-side route from the view-family architecture to the cluster-graph equality case, while meet-witnessing and fiber coherence are stronger sufficient conditions. Under these conditions, the asymptotic Shannon capacity and the fixed Lovász-$\vartheta$ upper collapse to the logarithm of the number of connected components or realized transcript fibers, depending on the structure available. (L: MFT85, MFT91, MFT96, MFT102, MFT109)
@@
Finite converse foundation. Let $X$ be a finite latent source, let $Y$ be the deterministic observation transcript available to a resolver, and let $T$ be a finite auxiliary tag. Exact zero-error recovery from $(Y,T)$ is possible exactly when the pair map $x\mapsto(Y(x),T(x))$ is injective on the surviving ambiguity class. The same obstruction appears as confusability, counting, conditional-entropy, decoder-output, and finite-gap formulations, and it admits a budgeted finite-error extension. In particular, exact $k$-way resolution requires at least $\log_{2}k$ bits of side information.(L: OBS1-2, OBS5, CIA3, PRB47, PRB55, PRB68, PRB81, PRB83, PRB85, PRB87, PRB90)
@@
Boundary corollary. For deterministic multi-location encodings, the zero-incoherence threshold equals $1$: the single-source regime is the unique no-failure corner, and it is exactly the structural-integrity regime. In coding-theoretic terms, this is the erasure-correcting regime where the syndrome uniquely determines the message. The richer mathematics begins once the finite obstruction survives and acquires nontrivial confusability structure.(L: COH1-2, RAT1, CAP2-3)
@@
Secondary fact-side question. The same framework also supports a second structural question: which fact coordinates are determined by others across the realized state family? Under an affine restriction, that coordinate-dependence question becomes the standard span-membership condition on coordinate functionals, so the induced fact-closure is a representable matroid on fact indices. The value of this specialization here is operational rather than linear-algebraic novelty: the same realized-state model then carries two complementary invariants, a graph on latent states and, in the affine regime, a matroid on coordinates whose rank gives upper bounds on confusability and capacity. Those bounds are algorithmically tractable once the affine family is presented explicitly by a basis or generator matrix for its direction space; under that representation, the relevant rank quantity is exactly the rank of the restricted coordinate projection and satisfies formal size bounds such as $t(S)\leq|S|$, so the computation reduces to Gaussian elimination.(L: AFM1-11)
I-B Realizability and Verification
@@
The same deterministic model also raises a realizability question at its opposite boundary: when can a concrete host system realize the rate-$1$ corner and thereby certify structural integrity? We show that two structural properties are required:
@@ I-C Paper Organization
The main theorem arc is: partial views induce a confusability graph; exact recovery becomes graph colorability; block composition turns that graph into strong powers; the resulting normalized rates converge to the classical Shannon-capacity value of the induced graph; a fixed Lovász-$\vartheta$ upper theory bounds that value; and transitivity of confusability gives the model-side export to the classical cluster-graph equality case. The deterministic converse is the finite foundation of that arc. The affine and realizability sections are later specializations and boundary consequences of the same model rather than separate primary claims.
Section III gives the minimum model and threshold setup. Section IV develops the finite converse foundation. Sections VVIII contain the main graph-capacity arc: graph characterization, block and asymptotic capacity, upper theory, and equality characterization. Section IX then records the affine fact-side specialization as a second lens on the same confusability structure. Section X, Section XI, and Section XII return to the unit-rate boundary of the same model and develop its structural, operational, and rate-complexity consequences. Appendix A records representative host-level readings, and Supplement A contains case-study details and traces. A supplementary Lean 4 artifact machine-checks the converse family, graph extension, theta upper theory, affine fact-matroid specialization, threshold chain, operational realizability criterion, and rate-complexity arguments [7].
@@ I-E Contributions
The main contributions are grouped into the core graph-capacity arc and the application consequences built on the same model.
Core graph-capacity contributions.
  • A multi-fact confusability-graph extension in which partial observations on latent tuples produce non-clique confusability graphs, exact recovery becomes equivalent to graph colorability, success-set exactness becomes equivalent to induced-subgraph colorability, and the exact finite weighted-success value is characterized by the maximum $T$-colorable induced subgraph.
  • A strong-power block law and asymptotic Shannon-capacity export in which the $n$-block confusability graph is the $n$-fold strong power of the one-shot graph, so the classical Shannon-capacity framework applies directly to the induced graph family and the normalized block-rate sequence converges to the corresponding asymptotic capacity value.
  • A fixed Lovász-$\vartheta$ upper theory and standard equivalent forms in which the same asymptotic capacity is bounded above by complement-chromatic and fixed Lovász-$\vartheta$ upper objects, with standard orthonormal, primal-PSD, and dual-theta forms.
  • A model-side equality characterization in which the classical cluster-graph equality case is exported back to the original view-family model under transitivity of confusability, with meet-witnessing as a checkable sufficient condition and fiber coherence as a stronger one.
  • A deterministic finite converse toolkit for exact consistency, presented through pair injectivity together with confusability, counting, conditional-entropy, decoder-output, and entropy-gap formulations of the same finite obstruction, deterministic data processing, and a budgeted finite-error extension.
Consequence contributions.
  • An affine fact-matroid specialization from the same latent-state framework in which affine realized-state families induce a representable matroid on fact indices: semantic determination becomes span membership, and the resulting matroid rank yields tractable upper bounds for the main confusability/capacity problem.
  • Threshold and realizability corollaries showing that unit rate is the unique zero-incoherence regime, that unit rate is exactly the structural-integrity regime, that unit rate has $O(1)$ manual update cost while higher rates incur an $\Omega(n)$ lower bound, and that causal propagation together with provenance observability gives an operational criterion for realizable verifiable structural integrity in concrete hosts.
  • A supplementary machine-checked artifact verifying the theorem chain in Lean 4.
II Figures
[…]
III Model and Basic Quantities
@@
III-A Model assumptions and notation
@@
III-B Latent states, ambiguity, and independent rate
@@
Definition III.1 (Independent Location).
@@
Definition III.2 (Derived Location).
@@
Definition III.3 (Degrees of Freedom).
The degrees of freedom (DOF), or independent rate, of an encoding system for fact $F$ is the number of pairwise independent source locations encoding $F$ under the specified admissible edit set $E(C)$. In the single-fact model, the remaining encoding locations are treated as derived views of those independent sources.
Definition III.4 (Structural Integrity).
@@
Definition III.5 (Integrity Violation).
@@
Proposition III.6 (Unit Independent Rate Guarantees Coherence).
If the independent rate is $1$, then all reachable states are coherent.
(L: COH1)
@@
At independent rate $1$, exactly one location is independent and all remaining locations are derived from it. Derived locations cannot diverge from their source, so disagreement is unreachable.  $\blacksquare$
Proposition III.7 (Independent Rate Above One Permits Incoherence).
If the independent rate exceeds $1$, then incoherent states are reachable.
(L: COH2)
Proof.
With at least two independent locations, admissible edits can assign different values to those locations, producing a reachable disagreement state.  $\blacksquare$
III-C Zero-Incoherence Threshold
Definition III.8 (Zero-Incoherence Threshold).
Corollary III.9 (Threshold Corollary).
(L: RAT1, CAP2-3)
Proof.
Achievability. Suppose $\mathrm{DOF}(C,F)=1$, and let $L_{s}$ be the unique independent location. By Definition III.3, every other encoding location is non-independent and is therefore treated in this model as a derived view of $L_{s}$. By Definition III.2, updating $L_{s}$ determines the value of each such derived location without an additional manual edit. Hence all encoding locations agree in every reachable state, so zero incoherence is achieved.
Converse. Suppose $\mathrm{DOF}(C,F)=k>1$. By Definition III.1, there exist at least two independent locations $L_{1},L_{2}$ and an admissible edit sequence that can make them disagree. Equivalently, choose values $v_{1}\neq v_{2}$ and apply edits setting $L_{1}\leftarrow v_{1}$ and then $L_{2}\leftarrow v_{2}$. By independence, the second edit does not force $L_{1}$ to change. The resulting reachable state satisfies $\mathrm{value}(L_{1})\neq\mathrm{value}(L_{2})$ and is therefore incoherent.
@@ III-C Zero-Incoherence Threshold
Corollary III.10 (Exact Structural-Integrity Threshold).
A deterministic multi-location encoding has structural integrity if and only if its independent rate equals $1$.
(L: COH1-2)
Proof.
If the independent rate is $1$, Proposition III.6 shows that all reachable states are coherent, hence Definition III.4 holds. If the independent rate exceeds $1$, Proposition III.7 gives a reachable incoherent state, so structural integrity fails.  $\blacksquare$
This threshold is an early architectural consequence of the deterministic model. It identifies the unique trivial no-failure corner and, equivalently, the exact structural-integrity regime. Above that boundary, integrity violations are reachable by construction. The deeper theorems of the paper begin once one asks how much finite side information is needed whenever the transcript leaves a nontrivial ambiguity class.
Corollary III.11 (Side-Information Lower Bound).
If a surviving ambiguity class has size $k$, then exact zero-error resolution requires at least $\log_{2}k$ bits of side information.
(L: CIA1)
Proof.
Exact resolution of $k$ surviving alternatives requires at least $k$ distinguishable tags, hence at least $\log_{2}k$ bits.  $\blacksquare$
@@
IV Unified Finite Converse
This section supplies the one-shot obstruction that the later graph-capacity theory lifts from clique-shaped ambiguity to non-clique failure geometry. Its role is foundational and unifying rather than graph-theoretically novel: in the single-fact setting, every surviving ambiguity class is effectively clique-shaped under the observation transcript, so the same deterministic obstruction can be stated as injectivity, clique counting, entropy, decoder-output, and finite-error bounds. The cleanest zero-error formulation starts from the joint observation-tag map itself. The first theorem gives the basic criterion; the remaining statements package the same obstruction in the forms reused later in the graph and realizability arcs.
Theorem IV.1 (Pair-injectivity characterization).
Let $X$ range over a finite latent alphabet, let $Y$ be the deterministic observation transcript, and let $T$ be a finite auxiliary tag. Exact zero-error recovery from $(Y,T)$ is possible if and only if the pair map $x\mapsto(Y(x),T(x))$ is injective on the ambiguity class under consideration.
(L: OBS1-2)

### 31 Mar 2026 v4

# Exact Recovery Under Deterministic Partial Views: Confusability Graphs, Strong Powers, and Capacity

@@ Abstract
[…]
@@ I Introduction
The substantive question is what exact ambiguity structure survives when a finite latent source is observed only through deterministic partial views, and what zero-error recovery laws are forced by that structure. This paper answers that question with a graph-capacity theory for deterministic partial views on latent tuples. When applied to representative channel families underlying widely deployed infrastructure, the resulting realizability criterion proves that the majority of deterministic partial-view architectures operate outside the verifiable zero-error regime, rendering the preceding capacity limits operationally unavoidable.
The theory is architectural but information-theoretic: once the admissible view family is fixed, the induced confusability graph, exact tag budget, block law, asymptotic capacity, and standard upper bounds are all determined by that structure.
When a latent fact or tuple of facts is exposed only through partial observations, the observation need not identify a unique latent state. The surviving ambiguity is not merely a cardinality obstruction: the admissible view family induces a confusability graph on latent tuples. Its edges record the exact state pairs that the architecture cannot separate, proper colorings quantify the side-information budget needed for exact recovery, and repeated composition lifts the same structure to strong powers and asymptotic Shannon capacity. The resulting theory lies at the intersection of zero-error information theory, side-information source coding, and graph-capacity theory [1, 2, 3, 4, 5, 6, 7, 8].
Novelty. Unlike Witsenhausen’s setting where the observation law is given, here a deterministic multi-location partial-view architecture generates that law. [^1] Lean: MFT125-131 The exact coordinate-view model realizes confusability relations determined by upward-closed families of coordinate-agreement sets, and conversely every upward-closed family is realizable.222: MFT125-128 The model includes non-clique graphs, the cluster-collapse equality subclass, and is closed under block-composition (strong product).333: MFT24, MFT120-124
@@ I-A Distributed Source Coding Formulation
Failure geometry from partial views. The main jump of the paper is from clique-shaped ambiguity to architecture-specific failure structure. In the multi-fact partial-observation extension, the confusability graph is the structural object controlling exact consistency under partial views, exact recovery is equivalent to graph colorability, and exact finite weighted success is determined by the maximum $T$-colorable induced subgraph. The binary two-fact square already shows that the induced graph need not be a clique: it is a $4$-cycle, so failures above the unit-rate boundary are structured rather than uniform. 444: MFT3, MFT5-10, MFT24
@@
Asymptotic rate and upper theory. This graph characterization then lifts to strong powers and asymptotic zero-error rates. The full $n$-block confusability graph is the $n$-fold strong power of the one-shot graph; the normalized block-rate sequence converges to a real asymptotic Shannon-capacity value equal to the supremum of the finite block rates; and that value is bounded above by both the logarithm of the complement chromatic number and a fixed Lovász-$\vartheta$ upper convention. The one-shot upper object matches standard orthonormal-representation, primal-PSD, and dual-theta forms. 555: MFT35, MFT48, MFT55, MFT64, MFT69, MFT84
@@
Equality characterization. The paper also proves an equality characterization. For the original clique-fiber subclass coming from a surjective label map, asymptotic Shannon capacity and the fixed Lovász-$\vartheta$ upper both collapse to $\log|\mathcal{B}|$, where $\mathcal{B}$ is the fiber-label alphabet. This equality for cluster-graph structure is classical; the new point here is that transitivity of base confusability gives a model-side route from the view-family architecture to the cluster-graph equality case, while meet-witnessing and fiber coherence are stronger sufficient conditions. Under these conditions, the asymptotic Shannon capacity and the fixed Lovász-$\vartheta$ upper collapse to the logarithm of the number of connected components or realized transcript fibers, depending on the structure available. 666: MFT85, MFT91, MFT96, MFT102, MFT109
@@
Finite converse foundation. Let $X$ be a finite latent source, let $Y$ be the deterministic observation transcript available to a resolver, and let $T$ be a finite auxiliary tag. Exact zero-error recovery from $(Y,T)$ is possible exactly when the pair map $x\mapsto(Y(x),T(x))$ is injective on the surviving ambiguity class. The same obstruction appears as confusability, counting, conditional-entropy, decoder-output, and finite-gap formulations, and it admits a budgeted finite-error extension. In particular, exact $k$-way resolution requires at least $\log_{2}k$ bits of side information. This converse family is the one-shot foundation on which the later graph-capacity theory is built.777: OBS1-2, OBS5, CIA3, PRB47, PRB55, PRB68, PRB81, PRB83, PRB85, PRB87, PRB90
@@
Boundary corollary. For deterministic multi-location encodings, the zero-incoherence threshold equals $1$: the single-source regime is the unique no-failure corner, and it is exactly the structural-integrity regime. In coding-theoretic terms, this is the erasure-correcting regime where the syndrome uniquely determines the message. The richer mathematics begins once the finite obstruction survives and acquires nontrivial confusability structure.888: COH1-2, RAT1, CAP2-3
@@
Secondary fact-side question. The same framework also supports a second structural question: which fact coordinates are determined by others across the realized state family? Under an affine restriction, that coordinate-dependence question becomes the standard span-membership condition on coordinate functionals, so the induced fact-closure is a representable matroid on fact indices. The value of this specialization here is operational rather than linear-algebraic novelty: the same realized-state model then carries two complementary invariants, a graph on latent states and, in the affine regime, a matroid on coordinates whose rank gives upper bounds on confusability and capacity. Those bounds are algorithmically tractable once the affine family is presented explicitly by a basis or generator matrix for its direction space; under that representation, the relevant rank quantity is exactly the rank of the restricted coordinate projection and satisfies formal size bounds such as $t(S)\leq|S|$, so the computation reduces to Gaussian elimination.999: AFM1-11
I-B Boundary Interpretation
@@
The same deterministic model also has a boundary interpretation at rate $1$: when can a concrete host system realize the unique no-ambiguity regime and certify that it has done so? We show that two structural properties are required:
@@ I-C Paper Organization
The main theorem family is: partial views induce a confusability graph; exact recovery becomes graph colorability; block composition turns that graph into strong powers; the resulting normalized rates converge to the classical Shannon-capacity value of the induced graph; a fixed Lovász-$\vartheta$ upper theory bounds that value; and transitivity of confusability gives the model-side export to the classical cluster-graph equality case. The deterministic converse is the finite foundation of that family. The affine and realizability sections are later specializations and boundary consequences of the same model rather than separate primary claims.
Section II gives the minimum model and threshold setup. Section III develops the finite converse foundation. Sections IVVII contain the main graph-capacity arc: graph characterization, block and asymptotic capacity, upper theory, and equality characterization. Section VIII records the affine specialization as a second lens on the same confusability structure. Sections IX, X, and XI then interpret the unit-rate boundary and its structural consequences. Appendix A records representative host-level classifications and companion case-study details. A supplementary Lean 4 artifact machine-checks the converse family, graph extension, theta upper theory, affine fact-matroid specialization, threshold chain, structural-consequence criterion, and rate-complexity arguments [9, 10].
@@ I-E Contributions
Core graph-capacity: Multi-fact confusability-graph extension where partial observations produce non-clique graphs, exact recovery equals graph colorability, success-set exactness equals induced-subgraph colorability. Strong-power block law: $n$-block confusability graph is the $n$-fold strong power of the one-shot graph. Fixed Lovász-$\vartheta$ upper theory. Model-side equality characterization via transitivity of confusability exporting cluster-graph equality. Deterministic finite converse toolkit (pair injectivity, confusability, counting, conditional-entropy, decoder-output, entropy-gap formulations).101010: MFT3, MFT35, MFT48, MFT55, MFT64, MFT69, MFT84-85, MFT91, MFT96, MFT102, MFT109
Boundary consequences: Affine restriction yields representable matroid on coordinates with rank bounds. Unit rate is the unique zero-incoherence regime with $O(1)$ manual update cost vs $\Omega(n)$ for higher rates. Causal propagation plus provenance observability gives verifiable structural integrity criterion. Zero-incoherence classification of representative channel families (programming languages, databases, dependency managers). Supplementary Lean 4 artifact.111111: AFM1-11, COH1-2, RAT1, CAP2-3
II Model and Basic Quantities
@@
II-A Model assumptions and notation
@@
II-B Latent states, ambiguity, and independent rate
@@
Definition II.1 (Independent Location).
@@
Definition II.2 (Derived Location).
@@
Definition II.3 (Degrees of Freedom).
The degrees of freedom (DOF) is used as shorthand for the independent rate, defined as the number of pairwise independent source locations encoding $F$ under the specified admissible edit set $E(C)$. In the single-fact model, the remaining encoding locations are treated as derived views of those independent sources.
Definition II.4 (Structural Integrity).
@@
Definition II.5 (Integrity Violation).
@@
II-C Zero-Incoherence Threshold
Remark. At independent rate $1$, all locations are derived from a single source, so disagreement is unreachable. At independent rate above $1$, admissible edits can assign different values to two independent locations, producing a reachable disagreement state.121212: COH1-2
Definition II.6 (Zero-Incoherence Threshold).
Corollary II.7footnotemark: footnotetext: : RAT1, CAP2-3 (Threshold Corollary).
@@ II-C Zero-Incoherence Threshold
Achievability. Suppose $\mathrm{DOF}(C,F)=1$, and let $L_{s}$ be the unique independent location. By Definition II.3, every other encoding location is non-independent and is therefore treated in this model as a derived view of $L_{s}$. By Definition II.2, updating $L_{s}$ determines the value of each such derived location without an additional manual edit. Hence all encoding locations agree in every reachable state, so zero incoherence is achieved.
Converse. Suppose $\mathrm{DOF}(C,F)=k>1$. By Definition II.1, there exist at least two independent locations $L_{1},L_{2}$ and an admissible edit sequence that can make them disagree. Equivalently, choose values $v_{1}\neq v_{2}$ and apply edits setting $L_{1}\leftarrow v_{1}$ and then $L_{2}\leftarrow v_{2}$. By independence, the second edit does not force $L_{1}$ to change. The resulting reachable state satisfies $\mathrm{value}(L_{1})\neq\mathrm{value}(L_{2})$ and is therefore incoherent.
@@
Remark. Equivalently, structural integrity (Definition II.4) holds if and only if the independent rate equals 11.141414: COH1-2
This threshold is an early architectural consequence of the deterministic model. It identifies the unique zero-error corner and, equivalently, the exact structural-integrity regime. Above that boundary, integrity violations are reachable by construction. The deeper theorems of the paper begin once one asks how much finite side information is needed whenever the transcript leaves a nontrivial ambiguity class.
The surviving ambiguity also forces an information-theoretic lower bound: exact resolution of a $k$-way class requires at least $\log_{2}k$ bits of side information (Theorem III.3 below).151515: CIA1
@@
III Unified Finite Converse
This section supplies the one-shot obstruction that the later graph-capacity theory lifts from clique-shaped ambiguity to non-clique failure geometry. Its role is foundational and unifying rather than graph-theoretically novel: in the single-fact setting, every surviving ambiguity class is effectively clique-shaped under the observation transcript, so the same deterministic obstruction can be stated as injectivity, clique counting, entropy, decoder-output, and finite-error bounds. The cleanest zero-error formulation starts from the joint observation-tag map itself.
Theorem III.1footnotemark: footnotetext: : OBS1-2 (Pair-injectivity characterization).
Let $X$ range over a finite latent alphabet, let $Y$ be the deterministic observation transcript, and let $T$ be a finite auxiliary tag. Zero-error recovery from $(Y,T)$ is possible if and only if the pair map $x\mapsto(Y(x),T(x))$ is injective on the ambiguity class under consideration.